OpenStudy (anonymous):

LINEAR FUNCTION HELP For a certain linear function, y = 17 when x = 8, and y = 26 when x = 11. What is the initial value of that function?

OpenStudy (osanseviero):

So you know that f(8) = 17 and f(11) = 26. You can write linear functions like y-y1 = m(x-x1) So your first step should be to calculate m (the slope)

OpenStudy (osanseviero):

When you get the slope, you just use one pair of coordinates and get the answer in the general form y= mx + b

OpenStudy (anonymous):

ok i got 3 as the slope. is that correct ?

OpenStudy (osanseviero):

Yes :)

OpenStudy (anonymous):

so can i set it up like 17=3(8)+b ?

OpenStudy (osanseviero):

Work with this first y-y1 = m(x-x1)

OpenStudy (osanseviero):

What do you get from this? (Use the one you want)

OpenStudy (anonymous):

im not sure if I'm doing this correct of not .... i got -9=3(-3)+b

OpenStudy (osanseviero):

Nope...You can't substitute into y and x because those changing. You can substitute the values of a point P(x1,y1)

OpenStudy (osanseviero):

So if you know that the slope is 10 and you know that the lines passes through (8,9), you can do: y-9= 10(x-8) y = 10x - 80 + 9 y = 10x -71

OpenStudy (anonymous):

oh ok so i don't plug them in for x and y, just how u have it here next to them

OpenStudy (anonymous):

y=3x-7 hows that ?

OpenStudy (osanseviero):

Yes. Think about the problem per se. You know that the values of x and y will be changing all the time. But you know the slope and a point, so you can draw all the function with that.

OpenStudy (osanseviero):

It is correct

OpenStudy (anonymous):

ok thank you so much i got it from here, thanks for your help i totally understand now

OpenStudy (osanseviero):

Good. You should still try to solve one more